OpenStudy (anonymous):

evaluate and explain how lim x approaches -4+ (abs(x+4)/(x+4))(x+2)

OpenStudy (anonymous):

\[\lim_{x\to -4^+}\] means \(x\) is approaching \(-4\) from numbers greater than \(-4\) if \(x>-4\) then \(|x+4|=x+4\)

OpenStudy (anonymous):

and also lim x approaches -4- (abs(x+4)/(x+4))(x+2)

OpenStudy (anonymous):

so you are looking at \[\frac{x+4}{(x+4)(x+2)}=\frac{1}{x+2}\]

OpenStudy (anonymous):

and then \[\lim_{x\to -4^+}\frac{1}{x+2}=\frac{1}{-4+2}=-\frac{1}{2}\]

OpenStudy (anonymous):

no the (x+2) is multiplied by the whole thing

OpenStudy (anonymous):

ok, that is not really the point anyway

OpenStudy (anonymous):

\[ |x+4| = \left\{\begin{array}{rcc} -x-4 & \text{if} & x <-4 \\ x+4& \text{if} & x \geq -4 \end{array} \right. \]

OpenStudy (anonymous):

ohh i understand now the absolute value is broken down to have 2 values

OpenStudy (anonymous):

and so \[\frac{|x+4|}{x+4}= \left\{\begin{array}{rcc} -1 & \text{if} & x <-4 \\ 1& \text{if} & x> -4 \end{array} \right.\]

OpenStudy (anonymous):

right, and when you divide by \(x+4\) you will either have \(1\) or \(-1\) depending on whether \(x>-4\) or \(x<-4\)

OpenStudy (anonymous):

ohh i understand makes sense thanks!!!!!

OpenStudy (anonymous):

yw